Fermi National Laboratory


Accelerator Update

Friday August 20
The day shift began with the Tevatron (TeV) ground fault search continuing, and with Operations monitoring the Antiproton Source (Pbar) stack.

Operations reported at 9:34 AM that Electrical Engineering (EE) technicians had disconnected various voltage taps from the main bus with the intention of conducting another hipot.

A Pbar expert began conducting Debuncher power supply studies at 10:58 AM.

Operations began turning on MiniBooNE and the Main Injector (MI) for studies at 11:52 PM. Switchyard (SY) will get turned on a little later.

Saturday August 21
The midnight shift began with Operations monitoring the Pbar stack,

Operations began sending slipstacking beam to the MI dump at 12:02 AM. They also established beam to MiniBooNE.

The MI expert ended his flattop orbit studies at 1:24 AM.

Operations established beam to the Accumulator at 1:29 AM and a Pbar expert began stacking studies.

Operations established beam to the Meson MTest experiment at 2:40 AM and to MCenter five minutes later.

Operations reported at 10:47 AM that EE techs were back in the TeV to continue their ground fault hunt.

At 11:01 AM, Operations reported that a Booster RF station (BRF16) power supply had tripped off for the first of three times in fifteen minutes. Operations bypassed the station.

EE support techs in the TeV reported at 11:47 AM that the ground fault had disappeared. TeV experts decided they would hook everything back up and turn on the TeV.

At 12:53 PM, Operations reported Pbar emittance problems. Operators adjusted the tunes and cooling to solve the problem.

Operations reported at 2:13 PM that they had started the TeV checkout.

At 3:22 PM, Operations reported that the TeV ground fault returned. However, after two minutes into the hipot the ground fault "burned off."

Pbar resumed stacking at 3:36 PM.

Operations began sending beam to MiniBooNE at 3:55 PM.

Operations reported at 4:21 PM that the TeV had completed a successful abort test.

A Recycler expert reported at 5:35 PM that he had completed his machine studies.

Operations put the TeV through a dry squeeze at 5:38 PM.

Operations established a 12+2x0 study-store at 6:45 PM. (A 12+2x0 study-store is like the 12x0 study-store but with two more bunches added at a different spot for an abort test.)

Operations put the TeV into shot setup at 7:24 PM.

At 9:02 PM, an expert flew a wire and reported that the proton bunches had high emittances. Operation had already begun unstacking the antiprotons so they continued and hoped for the best. A few minutes later operators noticed some loss of beam on the first shot, but with each subsequent transfer, the loss lessened.

Operations established store 3744 at 9:52 PM with and initial luminosity of 67.5E30. Pbar resumed stacking and a Recycler expert continued with his study.

At 11:45 PM, Operations reported that they had to turn off MIRF14 due to sparking. With the help of an expert, this station was eventually turned back on.

Sunday August 22
The midnight shift began with Operations monitoring store 3744, and with Pbar stacking.

Recycler studies completed at 1:10 AM.

Operations established beam to MCenter and MTest at 1:33 AM.

An MI expert began an aperture scan at 7:30 AM.

The D-Zero (D0) experiment reported at 9:22 AM that they were conducting a roman pot test.

A TeV expert reported at 10:12 AM that he had adjusted the chromaticity values so that during the next shot setup the dry squeeze could be skipped.

Operations reported at 10:12 AM that the Recycler was stashing antiprotons.

The MI expert left at 10:19 AM, but asked operators to continue making MI aperture scans.

At 11:15 AM, a MI south anode power supply tripped off due to a MI expert's study. He had loaded a few too many events into the Time Line Generator. Operations reported at 11:52 AM that a MI RF station (MIRF5) had tripped off with the anode power supply and operators couldn't reset the station. They called a RF expert for help.

At 12:07 PM, a Booster RF station (BRF18) tripped off and wouldn't reset. An RF expert began repairs.

At 12:35 PM, operators with the help of an expert reset MIRF5

Operations terminated store 3744 at 1:08 PM.

At 1:14 PM, MIRF7, 14, & 18 tripped off. They all reset okay.

Operations put the TeV into shot setup at 1:23 PM and at 3:22 PM established store 3745. Its initial luminosity was 62.78E30. Pbar resumed stacking and Operations resumed sending beam to MCenter.

A Preaccelerator expert switched from the I- Source to the H- Source at 5:15 PM.

At 5:19 PM, operators turned off MiniBooNE and prepared the area for access.

Operations reported at 5:25 PM that Linac and Booster experts were running beam from the H- Source to the Booster dump to conduct Preaccelerator notcher studies.

Operations reported that MI and Recycler were turned off at 5:32 PM.

Operations reported that SY was turned off at 6:18 PM.

Booster studies ended at 8:41 PM and operators turned off the machine.

Monday August 23
The midnight shift began with Operations monitoring store 3745, and that the MI, MiniBooNE, Recycler, Booster, and Fixed Target machines were all off in preparation for shutdown.

A Pbar expert completed his studies at 2:50 AM. Operators began preparing the Pbar rings for access.

Operations terminated store 3745 at 4:06 AM. Operators began preparing the TeV tunnel for access.

Plans for Monday
Fermilab has officially begun its 13-week shutdown for maintenance and upgrades. The Accelerator Updates may not be published as often at three times a week, but they will continue at least as a once per week effort dealing with reports on upgrades, machine status, and some basic information on how all the machines work and what people have to do to keep them working. If you have any questions about our particle accelerators send them to accelupdates@fnal.gov and we will do our best to answer them.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/23/2004   email Fermilab